Whirly Corporation’s contribution format income statement for the most recent month is shown below:

Total Per Unit
Sales (8,900 units) $ 284,800 $ 32.00
Variable expenses 169,100 19.00
Contribution margin 115,700 $ 13.00
Fixed expenses 54,300
Net operating income $ 61,400

Required:

(Consider each case independently):

1. What would be the revised net operating income per month if the sales volume increases by 100 units?

2. What would be the revised net operating income per month if the sales volume decreases by 100 units?

3. What would be the revised net operating income per month if the sales volume is 7,900 units?
Answer #1
Case 1 Case 2 Case 3
Unit Sales (8900+100)=9000 (8900-100)=8800 7900
Sales (9000*32)=288,000 (8800*32)=281600 (7900*32)=252800
Variable expenses (9000*19)=171,000 (8800*19)=167200 (7900*19)=150100
Contribution margin 117,000 114400 102700
Fixed expenses 54300 54300 54300
Revised net operating income $62700 $60100 $48400.